OpenStudy (anonymous):

what is the sum of the roots of the equation 2x^2+6x-7=0

OpenStudy (anonymous):

b/2a?

OpenStudy (anonymous):

think about the quadratic, one side is +sqrt(b^2 - 4ac), other side is minus added together you get 1, leaving -b/2a

OpenStudy (anonymous):

no, -b/a

OpenStudy (anonymous):

forgot to add the 2 -b's

OpenStudy (anonymous):

First, 2x^2 means 2*?*2 Next, 6x-7 means 6*?-7 Last, once you find whats missing add both sums then it will give you 0. http://www.virtualnerd.com/ Search your topic on this website. The lady explains step by step in each process.

OpenStudy (anonymous):

If you copy what she does on paper you will figure out what the missing numbers are

OpenStudy (anonymous):

k(x-a)(x-b) => k x^2-k(a+b)x+kab. k=2 -k(a+b)=6 a+b=-3 sum of roots=-3

OpenStudy (anonymous):

I think my way is easier :P